# Wormholes in Einstein-Dirac-Maxwell theory with identical spacetime asymptotics

### v1

@@ III.1 Asymptotic behavior and boundary conditions
We seek regular finite-energy solutions of five ordinary differential equations (10)-(14). Even before numerical solution of these equations, it is possible to estimate an asymptotic behavior of the solutions, bearing in mind that the spinor fields $\bar{u}$ and $\bar{v}$ are functions, which diminish exponentially as $x\to\pm\infty$. It is seen from the form of Eqs. (12) and (13) that, due to the presence of the term $\left(x^{2}+x_{0}^{2}\right)^{-1/2}$, they are not $Z_{2}$-symmetric. This implies that we should seek solutions that are asymmetric with respect to the origin of coordinates $x=0$. Then the corresponding asymptotic behavior of the spinor fields has the form
@@
Here $\bar{M}_{+}$ corresponds to the ADM (Arnowitt-Deser-Misner) mass of the configurations under consideration as measured by a distant observer when $x\to+\infty$ and $\bar{M}_{-}$ is the mass as measured when $x\to$-$\infty$; $\alpha$ is a constant.
@@ III.3 Mass and the circumferential radius
$\bar{M}_{\pm}\equiv\mu M_{\pm}/M_{p}^{2}=\pm\frac{1}{2}\lim_{x\to\pm\infty}x^{2}\partial_{x}e^{A}=\pm\frac{c_{k}}{8}\lim_{\bar{x}\to\pm 1}\frac{\partial_{\bar{x}}A}{1-\bar{x}^{2}},$ (23)
where $M_{p}$ is the Planck mass and the last expression in the above equation represents the mass in terms of the compactified coordinate $\bar{x}$ from Eq. (21).
@@ III.4.1 The case of $\bar{e}=0$
As $\bar{\Omega}$ increases, the masses grow smoothly, and in the limit $\bar{\Omega}\to-0$ they demonstrate a fast increase according to the law $\bar{M}_{\pm}\sim|\bar{\Omega}|^{-1}$. In this limit: (i) the masses $\bar{M}_{+}$ and $\bar{M}_{-}$ become equal; (ii) the metric function $g_{tt}\equiv e^{A}\to 0$, the metric function $B\approx 1$, and the dimensionless Kretschmann scalar $\bar{K}\equiv K/\mu^{4}=\bar{R}_{\alpha\beta\mu\nu}\bar{R}^{\alpha\beta\mu\nu}$ is practically equal to 0 over all of space; and (iii) for all the solutions, the ratio $\bar{Q}/\bar{M}_{\pm}\to 1$ from above.
@@ III.4.2 The case of $\bar{e}\neq 0$
Figure 4: The ratio of the charges of the configurations to the radius of their throat $\bar{Q}_{\pm}/\bar{R}_{\text{th}}$ as a function of $\bar{\Omega}$ for different values of $\bar{e}$. For the systems with $\bar{e}=0$, the charges $\bar{Q}_{+}$ and $\bar{Q}_{-}$ are equal, $\bar{Q}_{+}=\bar{Q}_{-}=\bar{Q}$. For the systems with $\bar{e}\neq 0$, there are turning points located at $\bar{\Omega}=\bar{\Omega}_{\text{crit}}$ where $\bar{Q}_{\pm}/\bar{R}_{\text{th}}\to 1$: the upper parts of the curves correspond to the charge $\bar{Q}_{+}$, and the lower parts – to the charge $\bar{Q}_{-}$.
@@
The results of numerical calculations are exemplified by Fig. 5 where typical spatial distributions of the field functions for the systems with three different $\bar{e}$ are shown. This figure demonstrates characteristic changes in the distribution of the spinor fields when, as the coupling constant $\bar{e}$ increases, their maxima/minima shift to the left. This results in the fact that the location of the throat (minimum of the metric function $\bar{R}$) shifts to the left as well, and such systems have essentially different values of the masses $\bar{M}_{+}$ and $\bar{M}_{-}$ (see Fig. 3). Notice also that although for any value of $\bar{e}$, as the total mass of the systems under consideration increases, in the vicinity of the throat, the metric function $g_{tt}\equiv e^{A}\to 0$ and the metric function $g_{rr}\equiv Be^{-A}$ demonstrates a fast growth, the Kretschmann scalar $\bar{K}\to 0$ over all of space.
@@ III.4.4 Approximate solutions
Note that for all the configurations considered in Sec. III.4.1 and III.4.2, it is typical that, as $\bar{\Omega}\to\bar{\Omega}_{\text{crit}}\approx-\bar{e}$, over all of space, the spinor fields $\bar{u}$ and $\bar{v}$ are approximately equal to 0, and the metric function $B\approx 1$. Accordingly, for such systems, the approximate solutions (27) are already valid over all of space, including the region in the vicinity of the throat. In this case $\bar{Q}_{\pm}\to\bar{M}_{\pm}$; this corresponds to the extremal Reissner-Nordström solution.
@@ III.4.5 Stability and the mass-radius relation
Here the total particle number $N_{f}$ can be associated with the Noether charge of the spinor field $Q_{N}$, which becomes an integer after quantisation, $Q_{N}=N_{f}$. The Noether charge is defined via the timelike component of the four-current $j^{\alpha}=\bar{\psi}\gamma^{\alpha}\psi$ as $Q_{N}=\int\sqrt{-g}j^{t}d^{3}x,$ where in our case $\sqrt{-g}j^{t}=B^{3/2}e^{-3A/2}\left(r^{2}+r_{0}^{2}\right)\sin{\theta}\left(\psi^{\dagger}\psi\right)$. For symmetric systems, the lower limit in the above integral is the point $r=0$ where a throat or an equator are located.
@@
$R_{\star\pm}=\frac{\int_{x_{\text{max}}}^{\pm\infty}\sqrt{-g}j^{t}R(r)dr}{\int_{x_{\text{max}}}^{\pm\infty}\sqrt{-g}j^{t}dr},$ (30)
@@
Figure 6: Left panel: the dimensionless effective radius $\bar{R}_{\star}\equiv\mu R_{\star}$ as a function of the parameter $\bar{\Omega}$ for the configurations with different values of $\bar{e}$. Right panel: the mass-radius relations for the configurations with different values of $\bar{e}$. The dashed lines mark the parts of the mass-radius curves where the binding energy is negative (energetically unstable systems).
@@ IV Conclusion
    It is typical of the configurations from the items (i) and (ii) that, as $\bar{\Omega}\to\bar{\Omega}_{\text{crit}}\approx-\bar{e}$, in all of space the spinor fields $\bar{u}$ and $\bar{v}$ are approximately equal to zero, and the metric function $B\approx 1$. For such systems, it is possible to find approximate analytical solutions in the form (27), which are already valid over all of space, including the region near the throat. These solutions correspond to the Reissner-Nordström solution, and when the charges of the system $\bar{Q}_{\pm}\to\bar{M}_{\pm}$, we deal with the extremal Reissner-Nordström solution.
@@
    Unlike the systems with nontrivial spacetime topology supported by a ghost scalar field and spinor fields [25, 26], all the configurations considered in the present paper have only one throat.

### v2

@@ III.1 Asymptotic behavior and boundary conditions
We seek regular finite-energy solutions of five ordinary differential equations (10)-(14). Even before numerical solution of these equations, it is possible to estimate an asymptotic behavior of the solutions, bearing in mind that the spinor fields $\bar{u}$ and $\bar{v}$ are functions, which diminish exponentially as $x\to\pm\infty$. It is seen from the form of Eqs. (12) and (13) that, due to the presence of the term $\left(x^{2}+x_{0}^{2}\right)^{-1/2}$, they are not $Z_{2}$-symmetric. (Notice that such an asymmetry is inherent not only to the wormhole solutions considered here but also to the domain wall solutions supported by a spinor field [19].) This implies that we should seek solutions that are asymmetric with respect to the origin of coordinates $x=0$. Then the corresponding asymptotic behavior of the spinor fields has the form
@@
Here the dimensionless $\bar{M}_{\pm}\equiv\mu M_{\pm}/M_{p}^{2}$ with $M_{p}$ being the Planck mass, and $\bar{M}_{+}$ corresponds to the ADM (Arnowitt-Deser-Misner) mass of the configurations under consideration as measured by a distant observer when $x\to+\infty$ and $\bar{M}_{-}$ is the mass as measured when $x\to$-$\infty$; $\alpha$ is a constant.
@@ III.3 Mass and the circumferential radius
$\bar{M}_{\pm}\equiv\mu M_{\pm}/M_{p}^{2}=\pm\frac{1}{2}\lim_{x\to\pm\infty}x^{2}\partial_{x}A=\pm\frac{c_{k}}{8}\lim_{\bar{x}\to\pm 1}\frac{\partial_{\bar{x}}A}{1-\bar{x}^{2}},$ (23)
where the last expression in the above equation represents the mass in terms of the compactified coordinate $\bar{x}$ from Eq. (21).
@@ III.4.1 The case of $\bar{e}=0$
As $\bar{\Omega}$ increases, the masses grow smoothly, and in the limit $\bar{\Omega}\to-0$ they demonstrate a fast increase according to the law $\bar{M}_{\pm}\sim|\bar{\Omega}|^{-1}$. In this limit: (i) the masses $\bar{M}_{+}$ and $\bar{M}_{-}$ become equal; (ii) the metric function $g_{tt}\equiv e^{A}\to 0$, the metric function $B\approx 1$, and the dimensionless Kretschmann scalar $\bar{K}\equiv K/\mu^{4}=\bar{R}_{\alpha\beta\mu\nu}\bar{R}^{\alpha\beta\mu\nu}$ is practically equal to 0 throughout all of space; and (iii) for all the solutions, the ratio $\bar{Q}/\bar{M}_{\pm}\to 1$ from above.
@@ III.4.2 The case of $\bar{e}\neq 0$
Figure 4: The ratio of the charges of the configurations to the radius of their throat $\bar{Q}_{\pm}/\bar{R}_{\text{th}}$ as a function of $\bar{\Omega}$ for different values of $\bar{e}$. For the systems with $\bar{e}=0$, the charges $\bar{Q}_{+}$ and $\bar{Q}_{-}$ are equal, $\bar{Q}_{+}=\bar{Q}_{-}=\bar{Q}$. For the systems with $\bar{e}\neq 0$, there are turning points located at $\bar{\Omega}=\bar{\Omega}_{\text{crit}}$ where $\bar{Q}_{\pm}/\bar{R}_{\text{th}}\to 1$: the upper parts of the curves correspond to the charge $\bar{Q}_{+}$, while the lower parts are for the charge $\bar{Q}_{-}$.
@@
The results of numerical calculations are exemplified by Fig. 5 where typical spatial distributions of the field functions for the systems with three different $\bar{e}$ are shown. This figure demonstrates characteristic changes in the distribution of the spinor fields when, as the coupling constant $\bar{e}$ increases, their maxima/minima shift to the left. This results in the fact that the location of the throat (minimum of the metric function $\bar{R}$) shifts to the left as well, and such systems have essentially different values of the masses $\bar{M}_{+}$ and $\bar{M}_{-}$ (see Fig. 3). Notice also that although for any value of $\bar{e}$, as the total mass of the systems under consideration increases, in the vicinity of the throat, the metric function $g_{tt}\equiv e^{A}\to 0$ and the metric function $g_{rr}\equiv Be^{-A}$ demonstrates a fast growth, the Kretschmann scalar $\bar{K}\to 0$ throughout all of space.
@@ III.4.4 Approximate solutions
Note that for all the configurations considered in Sec. III.4.1 and III.4.2, it is typical that, as $\bar{\Omega}\to\bar{\Omega}_{\text{crit}}\approx-\bar{e}$, throughout all of space, the spinor fields $\bar{u}$ and $\bar{v}$ are approximately equal to 0, while the metric function $B\approx 1$. Accordingly, for such systems, the approximate solutions (27) are already valid throughout all of space, including the region in the vicinity of the throat. In this case $\bar{Q}_{\pm}\to\bar{M}_{\pm}$; this corresponds to the extremal Reissner-Nordström solution.
@@ III.4.5 Stability and the mass-radius relation
[…]
@@
[…]
@@
Figure 6: Left panel: the dimensionless effective radius $\bar{R}_{\star}\equiv\mu R_{\star}$ as a function of the parameter $\bar{\Omega}$ for the configurations with different values of $\bar{e}$. Right panel: the mass-radius relations for the configurations with different values of $\bar{e}$. The dashed lines mark the parts of the mass-radius curves where the binding energy is negative (energetically unstable systems). The solid black line represents the mass-radius curve for Einstein-Dirac-Maxwell systems with trivial spacetime topology.
@@
In turn, it is of interest to compare the mass-radius relations found here with those obtained for localized configurations with trivial spacetime topology. For this purpose, in the right panel of Fig. 6, we have shown the corresponding dependence for a charged Dirac star with trivial spacetime topology obtained within Einstein-Dirac-Maxwell theory (see Ref. [39]). It is seen from this graph that the behavior of the mass-radius curve for the system with trivial topology differs considerably (both qualitatively and quantitatively) from that of the systems with nontrivial topology and close value of the coupling constant $\bar{e}=0.6$ shown by the corresponding curves for $\bar{M}_{+}$ and $\bar{M}_{-}$.
@@ IV Conclusion
    It is typical of the configurations from the items (i) and (ii) that, as $\bar{\Omega}\to\bar{\Omega}_{\text{crit}}\approx-\bar{e}$, in all of space the spinor fields $\bar{u}$ and $\bar{v}$ are approximately equal to zero, while the metric function $B\approx 1$. For such systems, it is possible to find approximate analytical solutions in the form (27), which are already valid throughout all of space, including the region near the throat. These solutions correspond to the Reissner-Nordström solution, and when the charges of the system $\bar{Q}_{\pm}\to\bar{M}_{\pm}$, we deal with the extremal Reissner-Nordström solution.
@@
    Unlike the systems with nontrivial spacetime topology supported by a ghost scalar field and spinor fields [26, 27], all the configurations considered in the present paper have only one throat. Nevertheless, since we have considered only a restricted range of values of the throat parameter $x_{0}$, we cannot rule out the possibility that configurations with two or more throats may exist for other values of the input parameters of the system. However, this issue requires special studies.